a sugar solution was made by mixing 7 qt. of a 15% solution and 4 qt. of a 70% sugar solution. what is the concentration of the mixture?
Answer by lwsshak3(11628)

a sugar solution was made by mixing 7 qt. of a 15% solution and 4 qt. of a 70% sugar solution. what is the concentration of the mixture?
**
let x= concentration of the mixture
15%*7+70%*4=x*11
11x=1.05+2.8=3.85
x≈3.85/11=.35
concentration of the mixture=35%